INFRARED-ACTIVE VIBRON BANDS ASSOCIATED WITH RARE GAS SUBSTITUTIONAL IMPURITIES IN SOLID HYDROGEN

\underline{PAUL L. RASTON}, and DAVID T. ANDERSON, Department of Chemistry, University of Wyoming, Laramie, WY 82071-3838.

Solid para-hydrogen (pH_{2}) crystals doped with part per million concentrations of rare gas (Rg) atoms display a new zero phonon absorption feature which correlates with the pH_{2} pure vibrational Q_{1}(0) transition. This Rg induced Q_{1}(0) absorption has been studied at high resolution for Ne, Ar, Kr, and Xe doped pH_{2} crystals. The more polarizable the Rg atom, the more intense and red shifted is the induced pH_{2} Q_{1}(0) feature. The frequency and lineshape of the transition provides information on how localized the vibron is around the Rg impurity. Comparison of the experimental data with a recent theoretical model is very favourable.^{a} In addition, the Rg atom perturbs the S_{1}(0) pH_{2} transition resulting in peaks that show fine structure which is interpreted as a lifting in the m_{J} degeneracy of this J=2 upper state. Preliminary studies of Xe atom doped ortho-deuterium will also be presented and discussed.

aR. J. Hinde, J. Chem. Phys. 119, 6 (2003).